Substitution Method; 1. 3x-6y=4; 5x-5y=2 2. 2x+3y=3;4x-6y=4


1
Expert's answer
2022-03-14T19:56:50-0400

"1.\\
3x-6y=4\\5x-5y=2\\
3x=4+6y\\
x=\frac{4+6y}{3}\\
5\cdot\frac{4+6y}{3}-5y=2\\
20+30y-15y=6\\
15y=-14\\
y=-\frac{14}{15}\\
x=\frac{4+6\cdot\frac{-14}{15}}{3}=\frac{-\frac{24}{15}}{3}=-\frac{8}{15}\\
(-\frac{8}{15},-\frac{14}{15})"

"2.\\
2x+3y=3\\4x-6y=4\\
2x=3-3y\\
x=\frac{3-3y}{2}\\
4\cdot\frac{3-3y}{2}-6y=4\\
6-6y-6y=4\\
-12y=-2\\
y=\frac{2}{12}\\y=\frac{1}{6}\\
x=\frac{3-3\cdot\frac{1}{6}}{2}=\frac{\frac{5}{2}}{2}=\frac{5}{4}\\
(\frac{5}{4},\frac{1}{6})"